OpenStudy (anonymous):

A house-painting company charges 376 dollars plus 12 dollars per hour. another painting company charges 280 dollars plus 15 dollars per hour. a. how long is a job for which both companies will charge the same amount? b. what will that cost be?

OpenStudy (anonymous):

y=12x+376 y=15x+280 find the intersection, do you know how to do that?

OpenStudy (anonymous):

if you don't have a calculator, just solve by substitution 15x+280=12x+376 3x=96 x=32 so after 32 hours

OpenStudy (anonymous):

then put in 32 for x... 15(32)+280=$760.00

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

anytime